OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (4, -4) with 0° ≤ θ < 360°. (4 square root of 2, 45°), (-4 square root of 2, 225°) (4 square root of 2, 135°), (-4 square root of 2, 315°) (4 square root of 2, 225°), (-4 square root of 2, 45°) (4 square root of 2, 315°), (-4 square root of 2, 135°)

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

can u find that angle and length ?

ganeshie8 (ganeshie8):

length = \(\large \sqrt{4^2 + 4^2} = 4\sqrt{2}\)

ganeshie8 (ganeshie8):

that angle is 45 degrees, as the sides are equal

ganeshie8 (ganeshie8):